Two points A and B are maintained at a potential of 7 V and -4 V respectively. The work done in moving 50 electrons from A to B is

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Work done = e * V = 50 * 1.6 * 10^-19 * 11 = 8.8 * 10^-17 Joule
 
Hope it helps
 
In case of any difficulty, Please feel free to ask again